OpenStudy (anonymous):

Give the function that describes the sequence. X Y 1 18 2 23 3 28 4 33 A. 10x + 3 B. 3x + 1 C. 11x - 9 D. 5x + 13

OpenStudy (raden):

use trial error let D as the anwer y = 5x + 13 if x = 1 then y = 5*1 + 13 = 18 (true) if x = 2 then y = 5*2 + 13 = 23 (true) if x = 3 then y = 5*3 + 13 = 28 (true) if x = 4 then y = 5*4 + 13 = 33 (true) because all be true so, dont bother it is D